OpenStudy (anonymous):

slope 3 starting at point x=3 y=6

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-6=3(x-3)\] \[y-6=3x-9\] \[y=3x-3\]